# Fix-Con: Automatic Fault Localization and Repair of Deep Learning Model Conversions between Frameworks

### v3

@@ IV-A1 Preprocessing (PP)
For some instances, differences in preprocessing settings between Source and Target can result in output label differences. We explore the effect of different preprocessing configurations for Targetusing the default preprocessing setting from the official repositories of the DL frameworks both Source and Target. For example, in the case of converting MobileNetV2 from PyTorch to Keras, Fix-Con examines the performance of the converted Target, using both official PyTorch and Keras preprocessing settings for MobileNetV2.
IV-A2 Input Dimensions (ID)
@@
Islam et al. [60] identified that changes in layer input dimensions can affect the performance of a DNN model. We also empirically identified that model conversion tools accept misconfigurations of Target model input dimensions without notifying the user of the mismatch between Source and Target input dimensions that may potentially result in output label discrepancies. We explore problems arising from such misconfigurations in this fault type. For example, we inspected the conversion of ResNet101 from PyTorch to TFLite, modifying the input dimensions from (1, 3, 224, 224) used in Source to (1, 3, 299, 299) used in Target. The conversion process completed without errors and warnings. However, the model presented a significant output label discrepancy between Source and Target (37373737%). In our fault localization approach, we check if input dimensions between Source and Target match, and report any mismatches to the fault repair component.
IV-A3 Tensor Shape and Structure (TSS)
@@
Another potential error identified by Islam et al. [60] relates to tensor shapes and changes in model structure. We consider cases where an erroneous tensor shape can affect the correctness of the model conversion where Source and Target models have differences in output labels. Such a scenario may be observed when a transpose operation is introduced by a converter tool to modify the model input. This is typically manifested as a Transpose node right after the input node in the computation graph. In an erroneous conversion scenario, the Transpose node might generate a tensor with correct shape, but incorrect structure. To illustrate this, consider the example of a Source input with the shape (1, 224, 224, 3). The Target model on the other hand has a shape of (1, 3, 224, 224) after applying a Transpose node that modifies the input of Target, by either applying Transpose(tensor, [0, 2, 3, 1]), or Transpose(tensor, [0, 3, 2, 1]), depending on the implementation. The modified shape is used in subsequent nodes in the computation graph. Owing to the ambiguity in dimensions where both transpose operations result in a tensor with correct shape expected by the subsequent layers in the Target model, it can result in model conversion errors as one of them has an incorrect structure. To localize faults like this, Fix-Con checks if there is a difference in model input shape between Source and Target. For cases presenting differences, Fix-Con further examines if any of the input-related subsequent layers contains a shape transformation operation (like Transpose) and considers such scenarios as potential sources of faults, conservatively. Potential faults related to tensor structure may also happen when a layer tensor in the Target model is incorrectly transformed by the converter and is followed by a flattening or a reshaping operation that loses the original structure information, making the difference with Source model non-obvious. In particular, if a tensor has a different shape between Source and Target (i.e., due to erroneous conversion) but the total number of elements is the same across tensors, a Flatten or a Reshape operation on such a tensor will make it impossible to detect the mismatch between Source and Target, as the structure information is lost. However, the order of elements between the corresponding Source and Target tensors will be different after the Flatten or a Reshape operation, which may result in output label discrepancies. For instance, if a layer has shape (1, 3, 32, 32) on Source  but was converted to (1, 32, 32, 3) on Target, and is followed by a Flatten node in both models, the result will be tensors with (3072) elements in both models. We localize this fault by examining the tensor shape of the dominator nodes preceding any Flatten or Reshape nodes, and check if it matches with the corresponding nodes in the Source model.
@@
For this category of faults, we examine and compare layer activations between Source and Target models. For layers where differences in activations are outside of an expected range of differences (as described in Section IV-A4), we mark those layers as suspicious and rank them so that we examine the most suspicious layer first for fault types CG, LH, and WB. We then iteratively proceed down this ranked list of suspicious layers for inspecting the three layer-based fault types, CG, LH and WB, discussed below. We start by describing our approach for producing a ranked list of suspicious layers in layer activation analysis followed by our fault localization approach for CG, LH, and WB with the ranked suspicious layers. An example instance of the differences between Source and Target in weights, biases, and hyperparameters that Fix-Con attempts to localize as potential faults is shown in Figure 3. We find, for a core computational layer under inspection, that the weights present small differences between Source and Target. Additionally, the pads hyperparameter that is present in Source, is missing on Target in Figure 3.
IV-A4 Layer Activation Analysis
@@
We compare the activations for each convolutional layer between Source and Target models, as a four-step process. (1) We extract two equally sized sets of images from the test dataset: one for similar images (images that produce the same top label prediction across Source and Target), and one for dissimilar images (images with different labels between Source and Target). We then extract the layer activation tensors of both sets by loading the Source and Target ONNX models to TVM and performing inference using the TVM Debugger [49]. The TVM debugger generates the complete model graph and parameters at compile time, while also gathering layer activations for each input. Using these activations and the Source ONNX model, we then utilize the LAYER_ANALYSIS function, shown in Algorithm 1. This function takes as input the original source ONNX model, as well as the activations for the sets of similar and dissimilar images (sim_acts, dis_acts). It also utilizes the List module, responsible for list length check and sorting operations, and the Kruskal-Wallis parametric test (KWTest). (2) To identify when activation mismatches between Source and Target indicate a potential problem, we first extract the differences in activations between Source and Target, per element within each layer, for the set of similar images and store these differences in a set, sim_diff in line $88$ of Algorithm 1. We treat differences in activations between Source and Target for these similar images as acceptable differences and we use it to estimate the expected difference in distribution of activations for correct conversions. (3) Next, for the set of dissimilar images, we compute the set of dissimilar differences, diss_diff in line $1010$, per tensor element per layer. We then check if the difference in activations per tensor element per layer belongs to the corresponding expected distribution of activation differences (from the previous step). To do this, we use the Kruskal-Wallis non-parametric statistical test (line $1111$) [57] [^1] The distribution of activation differences for correct conversion instances did not follow a normal distribution, so we chose the non-parametric Kruskal-Wallis test. to determine if sim_diff and diss_diff belong to the same distribution (at 5% significance level) (line $1515$). (4) From step 3, we obtain the tensor elements whose difference in activations for sim_diff and diss_diff did not belong to the same distribution, for each layer between Source and Target. These are considered as potentially problematic elements that may result in model conversion errors. We count the number of problematic tensor elements for each model layer (line $1616$) and rank the layers in descending order of number of problematic elements (line $1818$). This ranking of suspicious layers is the basis for fault localization and repair of fault types CG, LH, and WB. We only consider convolutional layers as these are the core computational layers for a Convolutional Neural Network (CNN). However, this can be modified to consider other layer types as well.
Algorithm 1 Layer Analysis in Fix-Con
@@
9:                 sim_diff $\leftarrow$ sim_diff $\cup$ sim[element]
10:             for diss in diss_acts do
11:                 diss_diff $\leftarrow$ diss_diff $\cup$ diss[element]
12:             kw_out $\leftarrow$ KWTest(sim_diff, diss_diff)
13:             // If the null hypothesis of
14:             // same distribution is rejected,
15:             // then the element counts as problematic.
16:             if kw_out.p_value <<< 0.05 then
17:                 layer.prob_elems+++++ +
18:    // Sort layers in descending order of problematic elements.
19:    return sort(model.layers,
              (l1, l2) $\Rightarrow$ l1.prob_elems -- l2.prob_elems)
IV-A5 Weights and Biases (WB)
@@
IV-A6 Layer Hyperparameters (LH)
Incorrectly converted hyperparameters are another potential source of error. For example, with a convolutional layer, we expect that the padding, strides, dilation, and other configurations to remain unchanged during model conversion. Starting from the most suspicious layer in the ranked list, we compare layer hyperparameters and attributes – (pads, strides, kernel_shape, dilations, epsilon, min, max, axis) between Source and Target. If we find any differences, we mark them as potential faults. In particular, we observe $33$ types of differences: (1) a hyperparamater is present in Source but not in Target; (2) a hyperparameter is present in Target but not in Source (e.g., added by the converter tool); and (3) a hyperparameter is present both in Source and Targetbut their values are different.
IV-A7 Computation Graph (CG)
@@
Figure 3: Indicative example of differences in layer weights and hyperparameters, introduced in the model conversion process.
IV-B Fault Repair
@@
IV-B1 Preprocessing Differences (PP) Repair
@@
IV-B2 Input Dimensions (ID) Repair
Following fault localization, if a difference is detected between Source and Target input dimensions, Fix-Con automatically fixes Target to match with the input dimensions from Source. We observed four erroneous model conversions in our experiment with incorrect input dimensions, with two of them requiring further actions in model nodes for tensor shape and structure (discussed below).
IV-B3 Tensor Shape and Structure (TSS) Repair
Following fault localization, as described in Section IV-A3, if the input is different between Target and Source in a problematic conversion, and input is succeeded by a Transpose node, then Fix-Con attempts to apply a fix by removing that node, in addition to adjusting the Target input to match Source. Furthermore, if Fix-Con detects any nodes that are related to input processing and normalization (e.g., InceptionV3 architecture contains Gather, Unsqueeze, Add, and Mul nodes before the added Transpose), then Fix-Con adjusts the attributes of these nodes in order to be compatible to the Source input dimension. In cases where Fix-Con detects the presence of Flatten or Reshape nodes that alter the shape of tensors, then it attempts to repair this by inserting a Transpose node right before Flatten or Reshape to match tensor shape with Source.
@@ IV-B3 Tensor Shape and Structure (TSS) Repair
IV-B4 Weights and Biases (WB) Repair
@@
If a mismatch is found between any weight or bias tensor across model layers by the fault localization component, then Fix-Con replaces the values from the corresponding tensor in Source to Target. As mentioned earlier, Fix-Con performs the repair starting from the most suspicious layer in the ranked list.
IV-B5 Layer Hyperparameters (LH) Repair
@@
Upon marking a layer hyperparamter case as potentially problematic, as described in Section IV-A6, Fix-Con attempts to fix Target by applying the corresponding setting of Source. In particular, if (1) a node is present in Source but is missing from the computation graph of Target, Fix-Con will update the corresponding node including the Source hyperparameter to Target. If (2) a hyperparameter is added in Target but is missing from Source, Fix-Con will remove it from the target node. And finally, if (3) there is a value mismatch between Source and Target for the hyperparameter, Fix-Con will update Target in accordance to the source value.
IV-B6 Computation Graph (CG) Repair
@@
In accordance to Section IV-A7, if a difference in subgraphs is detected between Source and Target, Fix-Con attempts to repair it by replacing the Target subgraph nodes with Source subgraph nodes, preserving all the node properties and ensuring all weights, biases and hyperparameters are correctly updated in the Target graph.
@@
Regarding the order in which our repair strategies are applied, we first applied the ones that were input-based, namely PP, ID, and TSS. We do PP first as changes to preprocessing may fix errors observed downstream. We then examine ID, with TSS being an additional restriction to ID mismatches. We then apply repairs that are “layer-based” – WB, LH, and CG. Among the ones with layer-based effects, we first perform repairs related to model weights, biases and hyperparameters, and then those related to computation graph modifications. Note that some graph transformations may also be related to hyperparameter changes, e.g., removing a Target Pad node to adapt to Source structure requires the addition of a padding hyperparameter to the next convolutional layer.
@@
We implemented Fix-Con in Python. We used Apache TVM Debugger for layer activations analysis. The debugger allows the extraction of model graph structure and parameters in metadata files, and a log of the layer activations throughout the model structure. In order to implement the repair strategies, we used the ONNX API to modify Target.
V Experiments
@@
We consider three widely used image recognition models of various sizes: MobileNetV2 [71], ResNet101V2 [58, 59], and InceptionV3 [72]. We use models pre-trained on ImageNet [55] using native model definitions and pre-trained parameters/weights sourced from four different DL frameworks’ repositories: Keras [54], PyTorch [67], TensorFlow(TF) [64], and TFLite [64]. For our experiments, we used two datasets: the ILSVRC2017 [70] object detection test dataset consisting of $55005500$ images, and the ImageNetV2 dataset [69] consisting of 10,0001000010,00010 , 000 images. We used the ILSVRC2017 dataset for detecting and repairing errors in model conversion. We then assessed the correctness of the repaired models on ImageNetV2, utilizing it as an independent dataset. We used the problematic conversion cases identified in [63] as our candidates for fault localization and repair. We focused on cases where the conversion process introduced label discrepancies between Source and Target, and there were $1515$ such erroneous conversions, as seen in Figure 1.
@@
RQ1:
How effective is Fix-Con at localizing faults in the $1515$ erroneous model conversions?
RQ2:
How effective is Fix-Con at repairing the faults in the $1515$ erroneous model conversions? We evaluate which of the repair strategies are used most frequently in model fixes.
To check the robustness of Fix-Con with respect to devices, we ran our experiments on two devices of varying capabilities, an Intel-based server featuring a high-end Nvidia Tesla K40c (GK11BGL) GPU, and a Laptop featuring a low-end Intel(R) GEN9 HD Graphics NEO. Our experiment results remained the same on both devices.
V-A Fault Analysis Setup

### v4

@@ 4.1.1. Preprocessing (PP)
For some instances, differences in preprocessing settings between Source and Target can result in output label differences. We explore the effect of different preprocessing configurations for Target using the default preprocessing setting from the official repositories of the DL frameworks both Source and Target. For example, in the case of converting MobileNetV2 from PyTorch to Keras, FetaFix examines the performance of the converted Target, using both official PyTorch and Keras preprocessing settings for MobileNetV2.
4.1.2. Input Dimensions (ID)
@@
Islam et al. (Islam et al., 2020) identified that changes in layer input dimensions can affect the performance of a DNN model. We also empirically identified that model conversion tools accept misconfigurations of Target model input dimensions without notifying the user of the mismatch between Source and Target input dimensions that may potentially result in output label discrepancies. We explore problems arising from such misconfigurations in this fault type. For example, we inspected the conversion of ResNet101 from PyTorch to TFLite, modifying the input dimensions from (1, 3, 224, 224) used in Source to (1, 3, 299, 299) used in Target. The conversion process completed without errors and warnings. However, the model presented a significant output label discrepancy between Source and Target (37373737%). In our fault localization approach, we check if input dimensions between Source and Target match, and report any mismatches to the fault repair component.
4.1.3. Tensor Shape and Structure (TSS)
@@
Another potential error identified by Islam et al. (Islam et al., 2020) relates to tensor shapes and changes in model structure. We consider cases where an erroneous tensor shape can affect the correctness of the model conversion where Source and Target models have differences in output labels. Such a scenario may be observed when a transpose operation is introduced by a converter tool to modify the model input. This is typically manifested as a Transpose node right after the input node in the computation graph. In an erroneous conversion scenario, the Transpose node might generate a tensor with correct shape, but incorrect structure. To illustrate this, consider the example of a Source input with the shape (1, 224, 224, 3). The Target model on the other hand has a shape of (1, 3, 224, 224) after applying a Transpose node that modifies the input of Target, by either applying Transpose(tensor, [0, 2, 3, 1]), or Transpose(tensor, [0, 3, 2, 1]), depending on the implementation. The modified shape is used in subsequent nodes in the computation graph. Owing to the ambiguity in dimensions where both transpose operations result in a tensor with correct shape expected by the subsequent layers in the Target model, it can result in model conversion errors as one of them has an incorrect structure. To localize faults like this, FetaFix checks if there is a difference in model input shape between Source and Target. For cases presenting differences, FetaFix further examines if any of the input-related subsequent layers contains a shape transformation operation (like Transpose) and considers such scenarios as potential sources of faults, conservatively.
Potential faults related to tensor structure may also happen when a layer tensor in the Target model is incorrectly transformed by the converter and is followed by a flattening or a reshaping operation that loses the original structure information, making the difference with Source model non-obvious. In particular, if a tensor has a different shape between Source and Target (i.e., due to erroneous conversion) but the total number of elements is the same across tensors, a Flatten or a Reshape operation on such a tensor will make it impossible to detect the mismatch between Source and Target, as the structure information is lost. However, the order of elements between the corresponding Source and Target tensors will be different after the Flatten or a Reshape operation, which may result in output label discrepancies. For instance, if a layer has shape (1, 3, 32, 32) on Source  but was converted to (1, 32, 32, 3) on Target, and is followed by a Flatten node in both models, the result will be tensors with (3072) elements in both models. We localize this fault by examining the tensor shape of the dominator nodes preceding any Flatten or Reshape nodes, and check if it matches with the corresponding nodes in the Source model.
@@
For this category of faults, we examine and compare layer activations between Source and Target models. For layers where differences in activations are outside of an expected range of differences (as described in Section 4.1.4), we mark those layers as suspicious and rank them so that we examine the most suspicious layer first for fault types CG, LH, and WB. We then iteratively proceed down this ranked list of suspicious layers for inspecting the three layer-based fault types, CG, LH and WB, discussed below. We start by describing our approach for producing a ranked list of suspicious layers in layer activation analysis followed by our fault localization approach for CG, LH, and WB with the ranked suspicious layers. An example instance of the differences between Source and Target in weights, biases, and hyperparameters that FetaFix attempts to localize as potential faults is shown in Figure 3. We find, for a core computational layer under inspection, that the weights present small differences between Source and Target. Additionally, the pads hyperparameter that is present in Source, is missing on Target in Figure 3.
4.1.4. Layer Activation Analysis
@@
We compare the activations for each convolutional layer between Source and Target models, as a four-step process. (1) We extract two equally sized sets of images from the test dataset: one for similar images (images that produce the same top label prediction across Source and Target), and one for dissimilar images (images with different labels between Source and Target). We then extract the layer activation tensors of both sets by loading the Source and Target ONNX models to TVM and performing inference using the TVM Debugger (tvm, 2023). The TVM debugger generates the complete model graph and parameters at compile time, while also gathering layer activations for each input. Using these activations and the Source ONNX model, we then utilize the LAYER_ANALYSIS function, shown in Algorithm 1. This function takes as input the original source ONNX model (used to iterate the matching layer activations for both Sourceand Target), as well as the activations for the sets of similar and dissimilar images for both Sourceand Target(contained in parameters sim_acts and diss_acts). It also utilizes the List module, responsible for list length check and sorting operations, and the Kruskal-Wallis parametric test (KWTest). (2) To identify when activation mismatches between Source and Target indicate a potential problem, we first extract the differences in activations between Source and Target, per element within each layer, for the set of similar images and store these differences in a set, sim_diff in lines $88$ and $99$ of Algorithm 1. We treat differences in activations between Source and Target for these similar images as acceptable differences and we use it to estimate the expected difference in distribution of activations for correct conversions. (3) Next, for the set of dissimilar images, we compute the set of dissimilar differences, diss_diff (lines $1111$ and $1212$), per tensor element per layer. We then check if the difference in activations per tensor element per layer belongs to the corresponding expected distribution of activation differences (from the previous step). To do this, we use the Kruskal-Wallis non-parametric statistical test (line 13131313(et al., 2014) [^2] The distribution of activation differences for correct conversion instances did not follow a normal distribution, so we chose the non-parametric Kruskal-Wallis test. to determine if sim_diff and diss_diff belong to the same distribution (using a standard 5% significance level) (line $1717$). (4) From step 3, we obtain the tensor elements whose difference in activations for sim_diff and diss_diff did not belong to the same distribution, for each layer between Source and Target. These are considered as potentially problematic elements that may result in model conversion errors. We count the number of problematic tensor elements for each model layer (line $1818$) and rank the layers in descending order of number of problematic elements (line $2020$). This ranking of suspicious layers is the basis for fault localization and repair of fault types CG, LH, and WB. We only consider convolutional layers as these are the core computational layers for a Convolutional Neural Network (CNN). However, this can be modified to consider other layer types as well.
Algorithm 1 Layer Analysis in FetaFix
@@
9:                 sim_diff $\leftarrow$ sim_diff $\cup$
10:                                (sim.tgt[element] - sim.src[element])
11:             for diss in diss_acts do
12:                 diss_diff $\leftarrow$ diss_diff $\cup$
13:                                (diss.tgt[element] - diss.src[element])
14:             kw_out $\leftarrow$ KWTest(sim_diff, diss_diff)
15:             // If the null hypothesis of
16:             // same distribution is rejected,
17:             // then the element counts as problematic.
18:             if kw_out.p_value <<< 0.05 then
19:                 layer.prob_elems+++++ +
20:    // Sort layers in descending order of problematic elements.
21:    return sort(model.layers,
22:                    (l1, l2) $\Rightarrow$ l1.prob_elems -- l2.prob_elems)
4.1.5. Weights and Biases (WB)
@@
4.1.6. Layer Hyperparameters (LH)
Incorrectly converted hyperparameters are another potential source of error. For example, with a convolutional layer, we expect that the padding, strides, dilation, and other configurations to remain unchanged during model conversion. Starting from the most suspicious layer in the ranked list, we compare layer hyperparameters and attributes – (pads, strides, kernel_shape, dilations, epsilon, min, max, axis) between Source and Target. If we find any differences, we mark them as potential faults. In particular, we observe $33$ types of differences: (1) a hyperparameter is present in Source but not in Target; (2) a hyperparameter is present in Target but not in Source (e.g., added by the converter tool); and (3) a hyperparameter is present both in Source and Targetbut their values are different.
4.1.7. Computation Graph (CG)
@@
Figure 3. Indicative example of differences in layer weights and hyperparameters, introduced in the model conversion process.
4.2. Fault Repair
@@
4.2.1. Preprocessing Differences (PP) Repair
@@
4.2.2. Input Dimensions (ID) Repair
Following fault localization, if a difference is detected between Source and Target input dimensions, FetaFix automatically fixes Target to match with the input dimensions from Source. We observed four erroneous model conversions in our experiment with incorrect input dimensions, with two of them requiring further actions in model nodes for tensor shape and structure (discussed below).
4.2.3. Tensor Shape and Structure (TSS) Repair
Following fault localization, as described in Section 4.1.3, if the input is different between Target and Source in a problematic conversion, and input is succeeded by a Transpose node, then FetaFix attempts to apply a fix by removing that node, in addition to adjusting the Target input to match Source. Furthermore, if FetaFix detects any nodes that are related to input processing and normalization (e.g., InceptionV3 architecture contains Gather, Unsqueeze, Add, and Mul nodes before the added Transpose), then FetaFix adjusts the attributes of these nodes in order to be compatible to the Source input dimension. In cases where FetaFix detects the presence of Flatten or Reshape nodes that alter the shape of tensors, then it attempts to repair this by inserting a Transpose node right before Flatten or Reshape to match tensor shape with Source.
@@ 4.2.3. Tensor Shape and Structure (TSS) Repair
4.2.4. Weights and Biases (WB) Repair
@@
If a mismatch is found between any weight or bias tensor across model layers by the fault localization component, then FetaFix replaces the values from the corresponding tensor in Source to Target. As mentioned earlier, FetaFix performs the repair starting from the most suspicious layer in the ranked list.
4.2.5. Layer Hyperparameters (LH) Repair
@@
Upon marking a layer hyperparameter case as potentially problematic, as described in Section 4.1.6, FetaFix attempts to fix Target by applying the corresponding setting of Source. In particular, if (1) a node is present in Source but is missing from the computation graph of Target, FetaFix will update the corresponding node including the Source hyperparameter to Target. If (2) a hyperparameter is added in Target but is missing from Source, FetaFix will remove it from the target node. And finally, if (3) there is a value mismatch between Source and Target for the hyperparameter, FetaFix will update Target in accordance to the source value.
4.2.6. Computation Graph (CG) Repair
@@
In accordance to Section 4.1.7, if a difference in subgraphs is detected between Source and Target, FetaFix attempts to repair it by replacing the Target subgraph nodes with Source subgraph nodes, preserving all the node properties and ensuring all weights, biases and hyperparameters are correctly updated in the Target graph.
@@
Regarding the order in which our repair strategies are applied, we first applied the ones that were input-based, namely PP, ID, and TSS, considering that issues introduced at this early processing stage can lead to problematic behavior accumulating and spreading throughout the model. We do PP first as changes to preprocessing may fix errors observed downstream. We then examine ID, with TSS being an additional restriction to ID mismatches. We then apply repairs that are “layer-based” – WB, LH, and CG. Among the ones with layer-based effects, we first perform repairs related to model weights, biases and hyperparameters, and then those related to computation graph modifications. Note that some graph transformations may also be related to hyperparameter changes, e.g., removing a Target Pad node to adapt to Source structure requires the addition of a padding hyperparameter to the next convolutional layer.
@@
We implemented FetaFix in Python. We used Apache TVM Debugger for layer activations analysis. The debugger allows the extraction of model graph structure and parameters in metadata files, and a log of the layer activations throughout the model structure. In order to implement the repair strategies, we used the ONNX API to modify Target.
5. Experiments
@@
We consider three widely used image recognition models of various sizes: MobileNetV2 (Sandler et al., 2018), ResNet101V2 (He et al., 2015, 2016), and InceptionV3 (Szegedy et al., 2015). We use models pre-trained on ImageNet (Deng et al., 2009) using native model definitions and pre-trained parameters/weights sourced from four different DL frameworks’ repositories: Keras (Chollet, 2015), PyTorch (Paszke et al., 2019), TensorFlow(TF) (Martín Abadi et al., 2015), and TFLite (Martín Abadi et al., 2015). For our experiments, we used two datasets: the ILSVRC2017 (Russakovsky et al., 2015) object detection test dataset consisting of $55005500$ images, and the ImageNetV2 dataset (Recht et al., 2019) consisting of 10,0001000010,00010 , 000 images. We used the ILSVRC2017 dataset for detecting and repairing errors in model conversion. We then assessed the correctness of the repaired models on ImageNetV2, utilizing it as an independent dataset, to verify our results and avoid dataset overfitting. We used the problematic conversion cases identified in (Louloudakis et al., 2023a) as our candidates for fault localization and repair. Both datasets contain real, non-synthetic samples and they are widely utilized by developers and researchers, while remaining tractable in size. We focused on cases where the conversion process introduced label discrepancies between Source and Target, and there were $1515$ such erroneous conversions, as seen in Figure 1. Overall, we selected models in the domain of computer vision for the following reasons: (1) To test our repair approach against the conversion faults we detected in the models we examined in our previous work (Louloudakis et al., 2023a); and (2) since our experiments involved 15 model conversion cases with numerous iterations and analysis within each, we chose image recognition models as they are more tractable and need less resources than other DL architectures, like transformers, for our experiments. However, our approach can generalize easily to any DL framework or conversion tool. It can be applied to any DL model architecture, including non-CNN models, only requiring the user to specify the core computational layers to be used in the comparison.
@@
RQ1::
How effective is FetaFix at localizing faults in the $1515$ erroneous model conversions?
RQ2::
How effective is FetaFix at repairing the faults in the $1515$ erroneous model conversions? We evaluate which of the repair strategies are used most frequently in model fixes.
To check the robustness of FetaFix with respect to devices, we ran our experiments on two devices of varying capabilities, an Intel-based server featuring a high-end Nvidia Tesla K40c (GK11BGL) GPU, and a Laptop featuring a low-end Intel(R) GEN9 HD Graphics NEO. Our experiment results remained the same on both devices.
5.1. Fault Analysis Setup